The following events apply to Tracey’s Restaurant for the Year 1 fiscal year:

  1. Started the company when it acquired $23,000 cash from the issue of common stock.
  2. Purchased a new cooktop that cost $22,000 cash.
  3. Earned $36,000 in cash revenue.
  4. Paid $20,000 cash for salaries expense.
  5. Paid $7,400 cash for operating expenses.
  6. Adjusted the records to reflect the use of the cooktop. The cooktop, purchased on January 1, Year 1, has an expected useful life of four years and an estimated salvage value of $3,500. Use straight-line depreciation. The adjusting entry was made as of December 31, Year 1.


Required
a. Record the events in general ledger accounts under an accounting equation. (Negative amounts should be indicated by a minus sign.)TRACEYS RESTAURANT Accounting Equation for Year 1 Assets Stockholders Equity Accum. Retained Cook Top = Com. Stock + Depr.

b. What amount of depreciation expense would Tracey’s report on the Year 2 income statement?

Depreciation expense

c. What amount of accumulated depreciation would Tracey’s report on the December 31, Year 2, balance sheet?

Accumulated depreciation
Answer #1

TRACEY'S RESTAURANT

Accounting Equation for Year 1

Assets = Stockholders Equity
Event Cash Cooktop Accumulated Depreciation = Common Stock + Retained Earnings
1. Issue Stock 23,000 = 23,000 +
2. Purchase Cooktop -22,000 22,000 = +
3. Revenue 36,000 = + 36,000
4. Paid Sal Exp -20,000 = + -20,000
5. Paid Op Exp. -7,400 = + -7,400
6. Depreciation Exp. - 4,625 = + - 4,625
= +
Totals 9,600 22,000 -4,625 = 23,000 + 3,975

Depreciation Expense for Each Year = Cost of Cooktop - Salvage Value / Useful Life in Years

Cost of Cooktop = $ 22,000

Salvage Value = $ 3,500

Useful Life in Years = 4 Years

Depreciation Expense for Each Year = 22,000 - 3,500 / 4

Depreciation Expense for Each Year = $ 4,625

Question B

Depreciation Expense for Year 2 = $ 4,625

Question C

Accumulated Depreciation December 31 Year 2 = 9,250

Accumulated Depreciation = Depreciation for Year 1 + Depreciation for Year 2

Accumulated Depreciation = 4,625 + 4,625 = $ 9,250